OpenStudy (anonymous):

f(x)= { (2x+p)/k for x=1,2,3,4,5 0 for all other values of x given that f(2) = 1/7 find p and k

hartnn (hartnn):

put x=0 in f(x) what u get ?

OpenStudy (anonymous):

zero...?

hartnn (hartnn):

yes, as f(0) = 0, u directly get p first

OpenStudy (anonymous):

p=0?

hartnn (hartnn):

f(x) =0 for any other value of x, so it will be 0 for x=0 yes! p=0 is correct :)

hartnn (hartnn):

now put x=2 in f(x)

OpenStudy (anonymous):

but the answer says p=1

hartnn (hartnn):

@sauravshakya could u verify the working ?

hartnn (hartnn):

is k = 28 given ?

OpenStudy (anonymous):

no~ k=35 >.<

hartnn (hartnn):

i wonder what mistake i did...

OpenStudy (anonymous):

f(6)=0 (2*6+p)/k=0 p=-12

OpenStudy (anonymous):

So, I think f(0)=0 cannot be done.

hartnn (hartnn):

lol. so u cannot find p like that way

OpenStudy (anonymous):

f(2)=1/7 (2*2+p)/k=1/7 28+7p=k

hartnn (hartnn):

thats the only equation....

OpenStudy (anonymous):

Yes.

hartnn (hartnn):

anything else is given ? @Kystal

OpenStudy (anonymous):

I think we are suppose to find the range of values of p and k

OpenStudy (anonymous):

The probability distribution function, f(x), for the discrete random variable X is given by the rule: f(x)= { (2x+p)/k for x=1,2,3,4,5 0 for all other values of x given that f(2) = 1/7 find p and k

hartnn (hartnn):

ohh!!

hartnn (hartnn):

now it can be solved

OpenStudy (anonymous):

THEN, f(1)+f(2)+f(3)+f(4)+f(5)=1

hartnn (hartnn):

^^ thats the 2nd equation

OpenStudy (anonymous):

Yep that will give another equation and u need to solve those two equations to find p and k

hartnn (hartnn):

can u find 2nd equation ?

OpenStudy (anonymous):

erm...30p/k=1 ?

hartnn (hartnn):

nopes, did u put 1,2,3,4,5 in place of x ? and equated to 1 ?

OpenStudy (anonymous):

yea...?

hartnn (hartnn):

(2+p+4+p+6+p+8+p+10+p)/k = 1 so the equation becomes ?

OpenStudy (anonymous):

ah 5p~!! miscount /.\

OpenStudy (anonymous):

got it!! thanks :)

hartnn (hartnn):

so could u find correct values of p,k ?

OpenStudy (anonymous):

p=1, k=35 XD

hartnn (hartnn):

:D